# Toward Robust Early Detection of Alzheimer’s Disease via an Integrated Multimodal Learning Approach

### v1

@@ 1 Introduction
Recently, deep neural network techniques have significantly advanced the classification of AD, overcoming the limitations of traditional manual diagnosis methods. Research in this field can be broadly divided into two categories: unimodal and multimodal classification tasks. Unimodal diagnosis often falls short in capturing the full complexity of AD, resulting in diagnostic challenges and high rates of misdiagnosis. Conversely, multimodal approaches integrate multiple types of data, including clinical information, cognitive scales, neuroimaging, and EEG signals, to provide a more comprehensive understanding of the disease. Cognitive scales, for example, offer a rapid and precise assessment of intellectual status and cognitive deficits [17]. Neuroimaging, particularly Magnetic Resonance Imaging (MRI), provides detailed images of brain structures, allowing for the observation of changes in brain volume and cortical atrophy. Electroencephalograms (EEG) objectively reflect brain neural activity, containing rich physiological and pathological information [18]. The integration of these multimodal data types enhances the accuracy and robustness of models, offering significant advantages over unimodal approaches.
@@ 3.1.1 Feature Tokenizer
where $b_{j}$ is the bias for the $j$-th feature. For numerical features, $f_{j}^{(\text{num})}$ is implemented as an element-wise multiplication with the vector $W^{(\text{num})}\in\mathbb{R}^{d}$. For categorical features, $f_{j}^{(\text{cat})}$ is implemented as a lookup table $W_{j}^{(\text{cat})}\in\mathbb{R}^{S_{j}\times d}$. The specific formulas are as follows:
@@ 4.1 Dataset and Experimental Details
Our experiments were conducted on our private ADMC dataset, which includes EEG, MRI, and scale data from 100 subjects. The subjects’ demographic details are as follows: mean age of 72.4 years, age range from 56 to 93 years, with 56 females and 22 married individuals. The dataset is divided into 80 samples for training and 20 samples for evaluation. We plan to make this private dataset publicly available to facilitate further research. We implemented the MSTNet network model using the PyTorch framework and trained it on an NVIDIA Tesla V100 GPU. During training, we set the dropout rate to 0.1, the batch size to 100, and the learning rate to 0.0001. The training process lasted for 24 hours.
@@ 4.2 Comparison Experiment
As presented in Table 1, we compare the proposed MSTNet model with several advanced unimodal and multimodal models. The unimodal models include HTCF [13], JD-CNN [11], and TableTransformer [10]. Among the multimodal models, we consider those developed by Zhang et al. [9], Qiu et al. [3], and Rad et al. [12]. The MSTNet model, equipped with the CMAA module, demonstrates superior performance over the unimodal models across all evaluated metrics. For instance, MSTNet effectively leverages spatial information from MRI and temporal information from EEG, outperforming TableTransformer in this regard. Furthermore, MSTNet solves existing multimodal classification models, which often face challenges in multimodal feature extraction and fusion. Compared to the best unimodal model, MSTNet improved its F1-score by 0.79%, and against the best multimodal model, MSTNet improved its F1-score by 4.17%.
@@ 4.3 Ablation Study
We conducted a series of ablation experiments to assess the impact of the DenseBlock module, TimesBlock module, CMAA module, and Feature Biases on the performance of MSTNet. The results are detailed in Table 1. The experiments reveal that the removal of each module results in varying degrees of performance decline. Specifically, the absence of the DenseBlock module severely hampers the model’s capability to process and extract intricate visual features from images, resulting in a 15% reduction in Accuracy. Similarly, removing the TimesBlock module significantly diminishes the model’s ability to capture temporal features from the EEG signal, leading to a 10% reduction in Accuracy. The exclusion of the CMAA module leads to a reduction in multimodal integration capabilities, resulting in a 15% reduction in Accuracy. Additionally, the model’s stability is compromised without the Feature Biases, causing a 5% reduction in Accuracy. These findings underscore the essential role of each module in maintaining the overall effectiveness of our proposed advanced MSTNet model.
@@ 5 Conclusion
In this study, we introduce an advanced multimodal classification model designed for the diagnosis of AD. This model overcomes the limitations of traditional single-modality approaches by effectively integrating diverse data types. The TimesBlock module is utilized to process EEG data, enabling the capture of complex temporal patterns. Additionally, a feature tagger and table encoding architecture are employed to enhance feature extraction from tabular data. Cross-modal attention aggregation is applied to integrate spatial MRI data with temporal EEG information. Our model has demonstrated superior performance in classifying AD, MCI, and NC, as confirmed by experiments conducted on our private ADMC dataset. By leveraging the strengths of multiple data modalities, our approach represents a significant advancement in the field of neurodegenerative disease classification.

### v2

@@ 1 Introduction
Recently, deep neural network techniques have significantly advanced the classification of AD, overcoming the limitations of traditional manual diagnosis methods. Research in this field can be broadly divided into two categories: unimodal and multimodal classification tasks. Unimodal diagnosis often falls short in capturing the full complexity of AD, resulting in diagnostic challenges and high rates of misdiagnosis. Conversely, multimodal approaches integrate multiple types of data, including clinical information, cognitive scales, neuroimaging, and EEG signals, to provide a more comprehensive understanding of the disease. Cognitive scales, for example, offer a rapid and precise assessment of intellectual status and cognitive deficits [17]. Neuroimaging, particularly Magnetic Resonance Imaging (MRI), provides detailed images of brain structures, allowing for the observation of changes in brain volume and cortical atrophy. Electroencephalograms (EEG) objectively reflect brain neural activity, containing rich physiological and pathological information [18]. The integration of these multimodal data types enhances the accuracy and robustness of models.
@@ 3.1.1 Feature Tokenizer
where $b_{j}$ is the bias for the $j$-th feature, and this process is called Feature Biases. For numerical features, $f_{j}^{(\text{num})}$ is implemented as an element-wise multiplication with the vector $W^{(\text{num})}\in\mathbb{R}^{d}$, and for categorical features, $f_{j}^{(\text{cat})}$ is implemented as a lookup table $W_{j}^{(\text{cat})}\in\mathbb{R}^{S_{j}\times d}$ as follows:
@@ 4.1 Dataset and Experimental Details
Our experiments were conducted on our private ADMC dataset, which comprises EEG, MRI, and scale data from 100 subjects. The subjects’ demographic details are as follows: mean age of 72.4 years, age range from 56 to 93 years, with 56 females and 22 married individuals. The dataset is divided into 80 samples for training and 20 samples for evaluation. We intend to release this dataset publicly to support additional research in the field. We implemented the MSTNet network model using the PyTorch framework and trained it on an NVIDIA Tesla V100 GPU. During training, we set the dropout rate to 0.1, the batch size to 100, and the learning rate to 0.0001. The training process lasted for 24 hours.
@@ 4.2 Comparison Experiment
As presented in Table 1, we compare the proposed MSTNet model with several advanced unimodal and multimodal models. The unimodal models include HTCF [13], JD-CNN [11], and TableTransformer [10]. Among the multimodal models, we consider those developed by Zhang et al. [9], Qiu et al. [3], and Rad et al. [12]. The MSTNet model, equipped with the CMAA module, demonstrates superior performance over the unimodal models in almost all metrics evaluated. Furthermore, MSTNet solves existing multimodal classification models, which often face challenges in multimodal feature extraction and fusion, demonstrating exceptional results.
@@ 4.3 Ablation Study
We conducted a series of ablation experiments to assess the impact of the DenseBlock module, TimesBlock module, CMAA module, and Feature Biases on the performance of MSTNet. The results are detailed in Table 1. The experiments reveal that the removal of each module results in varying degrees of performance decline. Specifically, the absence of the DenseBlock module severely hampers the model’s capability to process and extract intricate visual features from images. Similarly, removing the TimesBlock module significantly diminishes the model’s ability to capture temporal features from the EEG signal. The exclusion of the CMAA module leads to a reduction in multimodal integration capabilities. Additionally, the model’s stability is compromised without the Feature Biases. These findings underscore the essential role of each module in maintaining the overall effectiveness of our proposed advanced MSTNet model.
@@ 5 Conclusion
In this study, we introduce an advanced multimodal classification model designed for the diagnosis of AD. This model overcomes the limitations of traditional single-modality approaches by effectively integrating diverse data types. Our model has demonstrated superior performance in classifying AD, MCI, and NC, as confirmed by experiments conducted on our private ADMC dataset. This study reflects the potential of multimodal deep learning models in improving AD diagnosis and emphasizes the importance of integrating multiple data sources to gain a comprehensive understanding and prediction of the disease, representing a significant advancement in the field of neurodegenerative disease classification.
6 Acknowledgment
This work was supported in part by the National Key Research and Development Program of China (2023YFE0114900), the MoE Humanities and Social Sciences Project (24YJCZH225), Guangdong Basic and Applied Basic Research Foundation (No. 2022A1515110570), Guangxi Key R&D Project (No. AB24010167), Innovation Teams of Youth Innovation in Science, Technology of High Education Institutions of Shandong Province (No.2021KJ088), Shenzhen Science and Technology Program (No. KCXFZ20201221173008022), and Shenzhen Stability Science Program 2022 (2023SC0073).